# Behavioral and Topological Heterogeneities in Network Versions of Schelling’s Segregation Model

### v2

@@ 6.3.2   Mixing
7   Results
7.1   Segregation (Assortativity)
As anticipated, the observed mean final assortativity was higher for all Schelling-tolerance simulation runs. As in Xie & Zhou [3], heterogeneity of tolerances did, on their own, result in reduced assortativity. Consistent with Gandica, Gargiulo, & Carletti’s [4] observations, increasing topological heterogeneity alone did not produce results with reduced assortativity when $\epsilon\approx 0.39$, in fact, any increase in topological heterogeneity for Schelling-tolerance simulations resulted in a marked increase in realized assortativity. The combination of both dimensions of heterogeneity resulted in progressive reductions in assortativity. Xie & Zhou-tolerance simulations generated assortativity more slowly and resulted in reduced levels of segregation compared with Schelling-tolerance simulations at every level of densification. The lowest levels of assortativity observed occurred with Xie & Zhou [3] tolerances and 128 densifications. Results are illustrated in Figure 6.
Figure 6: Final graph assortativity for each group of simulations.
7.2   Organization
@@ 7.2.1   Hub Composition
Otherwise, the mean vacancy degree tended lower. For simulations with Xie & Zhou tolerances on densified graphs, on average, individuals in the group with the highest tolerance thresholds ($\geq 0.57$) had distinctly higher degree centralities than those in the group with the lowest tolerance thresholds ($\leq 0.07$). This difference became more pronounced as the number of densifications increased. These results are illustrated in Figures 8-9.
@@ 7.2.2   Paired Tolerances
Figure 10: Heatmaps show the log frequencies of tolerance-level pairs for dissimilar neighbors. Top row: 0 densifications; bottom row: 128 densifications; left column: initial distribution; right column: final distribution.
D $H(T)_{0}$ $H(T)_{F}$
0 7.651 5.842
32 7.646 6.155
64 7.654 6.138
96 7.636 6.163
128 7.642 6.254
Table 3: Initial and final Shannon entropies for the distribution of pairs’ tolerances. At each level, organization of tolerances is observed.
@@ 7.4   Tolerance Repels Intolerance
When both heterogeneous tolerances and heterogeneous topology are represented, the influence of tolerant nodes is amplified by densifications. A densified cluster populated by dissimilar, tolerant nodes acts as a strong repellant for intolerant nodes as tolerant pairs will directly share more neighbors. Second, a densified cluster occupied by similar nodes may continue to attract similar nodes but can no longer repel a dissimilar individual with a sufficiently large tolerance threshold. The introduction of relatively few highly tolerant dissimilar individuals to such a cluster will motivate catastrophic outmigration of intolerant individuals. These dynamics lead some densified clusters to become stores of tolerance and diversity. As the level of densification increases, the presence of tolerant nodes and densified clusters makes extremely homogeneous areas rarer. As this occurs, the most intolerant nodes can no longer find suitably homogenous candidate vacancies and become increasingly connected to tolerant, dissimilar nodes; see Figure 10. Thus, both the organization of tolerances and the organization of color-types in the final networks are reduced; see Table 3.
8   Conclusion
@@
In this study, we investigated the impact of representing multiple dimensions of heterogeneity in network models of residential segregation. Models which combine heterogeneity of tolerances with heterogeneity of topologies showed substantially different behavior than those which employed only one dimension of heterogeneity. Ordered migration and the resultant clusters of tolerance break down intolerant homogeny. This has bearing on the important question raised previously: to what extent can the collective dynamics of individual preferences lead to residential segregation? These results indicate that these collective dynamics may not contribute to residential segregation as much as previously thought if sufficient heterogeneity is represented. At a minimum, these results indicate the necessity of representing both heterogeneity of tolerances and heterogeneity of topologies in network models of residential segregation, as the omission of one or the other will result in the loss of these important model behaviors. It is notable that the dynamics which break down segregation along racial lines appear to create segregation in another way: the most tolerant individuals can become separated from the least tolerant individuals. This finding, paired with the results in Sayama & Yamanoi [14], may explain ongoing cultural fragmentation between urban and rural areas. The heterogeneity of tolerances and topology necessary for simultaneous maintenance of cultural diversity and social cohesion manifests segregation between tolerance levels themselves. This segregation ensures that opportunities for cultural diffusion between the intolerant subsets of population will be scarce, ensuring the maintenance of cultural diversity within the network. Here, rather than a fully connected network with weighted cultural diffusion rates, the network self-organizes such that the topology moderates cultural diffusion. This dynamic occurs even in the absence of agents’ preferences to occupy central locations on the network. We hypothesize that the addition of such a secondary preference would lead to a more dramatic level of segregation between tolerance levels since highly tolerant individuals who are not motivated by dissimilar neighbors would primarily be driven by this secondary preference. In such a scenario, highly tolerant individuals would be driven to dense areas and push intolerant individuals to sparse areas. Further work is required to elaborate our understanding of emergent segregation between tolerance levels.

### v3

@@ 6.2.1   Hub Composition
Otherwise, the mean vacancy degree tended lower. For simulations with Xie & Zhou tolerances on densified graphs, on average, individuals in the group with the highest tolerance thresholds ($\geq 0.57$) had distinctly higher degree centralities than those in the group with the lowest tolerance thresholds ($\leq 0.07$). This difference became more pronounced as the number of densifications increased. These results are illustrated in Figures 6-7. These results were robust to increases in excess housing and the presence of a minority population.
@@ 6.2.2   Paired Tolerances
Figure 8: Heatmaps show the log frequencies of tolerance-level pairs for dissimilar neighbors. Left, 0 densifications; right, 128 densifications.
This result was robust to the presence of a minority population, but an increase in excess housing increased the organization of tolerance pairs by enabling intolerant individuals to avoid contact. Figure 8 shows the organization of tolerance pairs for dissimilar neighbors at 128 densifications when the excess housing rate is 30%.
Figure 9: This heatmap shows the log frequencies of tolerance-level pairs for dissimilar neighbors when the excess housing rate is 30%.
7   Discussion
@@ 7.2   Tolerance Repels Intolerance
When both heterogeneous tolerances and heterogeneous topology are represented, the influence of tolerant nodes is amplified by densifications. A densified cluster populated by dissimilar, tolerant nodes acts as a strong repellant for intolerant nodes as tolerant pairs will directly share more neighbors. Second, a densified cluster occupied by similar nodes may continue to attract similar nodes but can no longer repel a dissimilar individual with a sufficiently large tolerance threshold. The introduction of relatively few highly tolerant dissimilar individuals to such a cluster will motivate catastrophic outmigration of intolerant individuals. These dynamics lead some densified clusters to become stores of tolerance and diversity. As the level of densification increases, the presence of tolerant nodes and densified clusters makes extremely homogeneous areas rarer. As this occurs, the most intolerant nodes can no longer find suitably homogenous candidate vacancies and increasingly remain dissatisfied and connected to tolerant, dissimilar nodes; see Figure 8. Thus, both the organization of tolerances and the organization of color-types in the final networks are reduced; see Table 1. It is important to note, however, that the presence of excess housing can moderate these dynamics by providing intolerant individuals with more opportunities to avoid dissimilar neighbors.
8   Conclusion and Future Work
@@
In this study, we investigated the impact of representing multiple dimensions of heterogeneity in network models of residential segregation. Models which combine heterogeneity of tolerances with heterogeneity of topologies showed substantially different behavior than those which employed only one dimension of heterogeneity. Ordered migration and the resultant clusters of tolerance break down intolerant homogeny. This has bearing on the important question raised previously: to what extent can the collective dynamics of individual preferences lead to residential segregation? These results indicate that these collective dynamics may not contribute to residential segregation as much as previously thought if sufficient heterogeneity is represented. At a minimum, these results indicate the necessity of representing both heterogeneity of tolerances and heterogeneity of topologies in network models of residential segregation, as the omission of one or the other will result in the loss of these important model behaviors.
It is notable that the dynamics which break down segregation along racial lines appear to create segregation in another way: the most tolerant individuals can become separated from the least tolerant individuals. This finding, paired with the results in Sayama & Yamanoi [19], may explain ongoing cultural fragmentation between urban and rural areas. The heterogeneity of tolerances and topology necessary for simultaneous maintenance of cultural diversity and social cohesion manifests segregation between tolerance levels themselves. This segregation ensures that opportunities for cultural diffusion between the intolerant subsets of population will be scarce thereby maintaining cultural diversity within the network. Here, rather than a fully connected network with weighted cultural diffusion rates, the network self-organizes such that the topology moderates cultural diffusion. This dynamic occurs even in the absence of agents’ preferences to occupy central locations on the network such as in Gambetta, Mauro, & Pappalardo’s [15] relevance model. We hypothesize that the addition of such a secondary preference would lead to a more dramatic level of segregation between tolerance levels since highly tolerant individuals who are not motivated by dissimilar neighbors would primarily be driven by this secondary preference. In such a scenario, highly tolerant individuals would be driven to dense areas and push intolerant individuals to sparse areas. Further work should investigate the influence of spatial relevance on the dynamics of segregation in networks with multiple dimensions of heterogeneity.
This work has implications for social and organizational cultural diffusion. The results suggest that the organization of tolerance levels in a network can be influenced by the network’s topology. As such, the organization of cultural diffusion in a network may be influenced by the network’s topology. This effect would be most pronounced in social and organizations systems where segregation along cultural lines is likely to occur. Future work should investigate the impact of network topology on cultural diffusion in social and organizational systems.
In addition, this work may hold importance for policymakers. The results suggest that heterogeneity of topology can either moderate or amplify segregation depending on the proportion and density of available excess housing. If policymakers aim to reduce segregation in communities, they may consider replacing low-density housing with high-density housing. Simply continuing to add both types of housing stock will continue to provide opportunities for tolerant individuals to self-organize in high-density areas while intolerant individuals continue to retreat to low-density areas. Given the variability and flexibility of zoning ordinances in the United States, this may never be a feasible policy option. A comparative analysis of real-world mixing patterns in high-density and low-density areas would be a valuable next step in this research. Another approach may examine the proportions of low- and high-density housing included in municipalities’ development plans. Emphasis on low-density housing may be a predicator of future segregation and urban-rural cultural fragmentation while underemphasis on low-density housing may predict a decline in cultural diversity.